an object is accelerating uniformly from 8.0 m/s to 16.0 m/s in 10 seconds. How far does the object travel in 10 seconds?
1
Expert's answer
2020-12-21T04:32:20-0500

"v=v_0+at\to a=\frac{v-v_0}{t}=\frac{16-8}{10}=0.8(m/s^2)"


"s=\frac{v^2-v_0^2}{2a}=\frac{16^2-8^2}{2\cdot0.8}=120(m)". Answer